# PartSDF: Part-Based Implicit Neural Representation for Composite 3D Shape Parametrization and Optimization

### v1

@@ 3.2 Training
Figure 3: Non-watertight part supervision strategy. (a) Part decomposition of a chair, (b) a 2D slice of its signed distances (red/blue) with parts highlighted in color, and (c) the specific regions of spaces in where each part is supervised.
Let us consider a dataset $\mathcal{D}$ where each shape $\mathcal{S}$ is decomposed in up to $P$ parts, given as the segmentation of its surface. To establish initial poses for each part, we fit simple primitives—cuboids or cylinders—to the segmented parts and use these primitives’ poses as the part poses. Therefore, each element of the dataset becomes a tuple $\mathcal{D}=\left\{\left(\mathcal{S},\left\{\mathcal{P}_{p}\right\},\left\{\mathbf{p}_{p}\right\}\right)_{i}\right\}$ of the shape, its parts, and their poses, where we write $\mathbf{p}_{p}=\left(\mathbf{q}_{p},\mathbf{t}_{p},\mathbf{s}_{p}\right)$.
@@
For watertight shapes $\mathcal{S}$, PartSDF’s decoder is then trained in an auto-decoding fashion (Park et al., 2019), where model parameters $\theta$ and part latent vectors $\mathbf{z}_{p}$ are optimized jointly. Note that we can directly use the parts poses from $\mathcal{D}$ during training. When a part is missing from a specific shape, we assign it the average pose of that part over the dataset, allowing the model to naturally learn to output $\text{SDF}>0$ for nonexistent parts, without having to predict an existence score, as in  (Petrov et al., 2023).
Some parts $\mathcal{P}_{p}$, however, may not be watertight as they are a subset of $\mathcal{S}$’s surface that may not be closed. Rather than computing an approximate SDF or editing the parts to make them watertight (Wu et al., 2020; Petrov et al., 2023), we adopt a different supervision strategy. Each part is supervised using the full shape’s SDF, but only in the region of space where each 3D point is closest to this part and not the other. Formally, a part $\mathcal{P}_{p}$ is supervised in $\left\{\mathbf{x}\in^{3}\,|\,p=\arg\min_{i}d_{i}(\mathbf{x})\right\}$, where $d_{i}(\mathbf{x})$ is the projection distance of $\mathbf{x}$ to part $i$, as depicted by Figure 3. To further prevent overlaps, we incorporate a non-intersection loss that maintains spatial separation between parts.
@@
with an additional index $p_{i}$ of the part closest to $\mathbf{x}_{i}$, re-using $\mathcal{X}$ as a slight abuse of notation to explicit that we compute these losses on the same 3D samples. $\hat{s}_{p_{i}}$ is the predicted part SDF at those points.
Finally, we introduce a non-intersection loss that pushes the SDF of parts to be positive at each position $\mathbf{x}_{i}$ where at least two parts have SDF$<0$. Using $\hat{\mathcal{X}}$ as the subset of such 3D points, the loss is written
@@
where $\hat{\mathbf{s}}_{i}$ is the vector of predicted part SDF at point $\mathbf{x}_{i}$ and $\mathbf{w}_{i}=\text{softmax}\left(\tilde{\mathbf{s}}_{i}\right),$ defining $\tilde{\mathbf{s}}_{i}$ as $\hat{\mathbf{s}}_{i}$ with all positive values replaced by $\infty$- ∞. These weights pushes predicted SDFs that are closer to 0, more strongly towards $>0$.
@@ 3.3 Part Encoding and Generation
Once PartSDF’s decoder is trained, it is frozen to be used for downstream tasks, optionally in conjunction with separate specialized networks. While this requires training two networks independently, it makes the training of each one easier. Furthermore, the decoder need only been trained once and used again and again for the different tasks.
@@
To reconstruct a shape from a given modality, e.g., point clouds or images, encoders can be trained to directly predict part latents and poses, to then be decoded with PartSDF: with input data $\mathcal{I}$ corresponding to the shape of our training data $\mathcal{D}$, encoders are trained to map this new modality to the part latents and poses of our pre-trained decoder. As an example, we show how to reconstruct unseen shapes for which part decomposition is unknown in Section 4.1. We do this by training a point cloud encoder to predict part latents and poses, which are then refined in an auto-decoding strategy and part-agnostic manner.
@@
For shape generation with a coherent part set, generative models can be trained to map random noise to the parts parametrization. Generative Adversarial Networks (GAN) (Goodfellow et al., 2014) have been used to generate shape latents, often dubbed latentGANs in related work (Achlioptas et al., 2018; Chen & Zhang, 2019) such as PQ-Net (Wu et al., 2020). Recently, diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) have gained a lot of traction for their generative performances, even in shape (Zhang et al., 2023) or part-based generation (Koo et al., 2023). Generating new shapes with PartSDF can then be achieved by training such generative models on the part poses and latents of our pre-trained encoder and, even if not shown here, could be extended to conditional generation on images or texts (Zhang et al., 2023; Koo et al., 2023).
@@ 4 Experiments
Table 1: Shape reconstruction. We compute the average Chamfer-Distance CD (×$\times 10^{4}$), Intersection over Union IoU (%), and Image Consistency IC between reconstructions and corresponding test shapes. We also report average per-part Intersection over Union pIoU (%) for part-based methods. Ours-1P uses a single part to reconstruct the full shape and Ours-PC uses a point cloud encoder to get initial part latents and poses that are refined in a part-agnostic manner.
@@
Ours 1.27 1.60 1.27 98.02 94.43 97.19 0.931 0.978 0.943 94.89 90.42 91.40
Ours-PC 1.28 1.73 1.30 97.95 83.38 97.04 0.930 0.966 0.941 94.25 78.61 87.42
In this section, we demonstrate how our method can be applied to various tasks and compare it to several part-based baselines. Further experimental details and additional results are available in Appendix B and F, respectively.
Datasets
Obtaining high-quality composite shape data is challenging as many detailed datasets tend to remain private. Public datasets, such as ShapeNet (Chang et al., 2015) and PartNet (Mo et al., 2019), often suffer from over-segmentation and inconsistencies in part definitions across samples, or provide only coarse semantic segmentations that lack the granularity needed for precise part manipulation. To remedy this, we use three curated datasets that offer consistent and well-defined part decompositions: (1) Car, a hand-processed subset of ShapeNet in which the wheels of the cars are separated from the main body; (2) Mixer, a set of liquid mixers made of a central helix within a tube and two attach points (Vasu et al., 2022); and (3) Chair, a cleaned subset of ShapeNet chairs, segmented with PartNet’s semantic labels and further divided into individual legs and arms. There are 1046, 1949, and 1332 shapes in each dataset with 5, 4, and 8 parts, respectively. We use $80\%$ of the data for training and the remaining $20\%$ as the unseen test set.
Baselines.
For part-based baselines, we have unfortunately found no publicly available code for modern architectures such as ProGRIP (Deng et al., 2022), SPAGHETTI (O. et al., 2022) and PASTA (Li et al., 2024). We therefore use DAE-Net (Chen et al., 2023) that learns deformable part template in an unsupervised manner, BAE-Net (Chen et al., 2019) that uses weak supervision with 8 labeled shapes, and PQ-Net (Wu et al., 2020), a fully-supervised approach that supports shape generation. As these part-based baselines are several years old, we also compare against the state-of-the-art 3DShape2VecSet, or 3DS2VS for short (Zhang et al., 2023). It is not ideal because, not being part-based, it is not suitable for engineering design. However, it gives an upper bound on the accuracy that can currently be obtained if one ignores the part decomposition, which typically allows for higher precision than when enforcing it.
Metrics.
Reconstruction accuracy is assessed using three different metrics: Chamfer-Distance (CD) for surface accuracy, Intersection over Union (IoU) for volume, and Image Consistency (IC) (Guillard et al., 2022) for shape appearance and normals. Part reconstruction is evaluated by averaging the per-part IoU and, as we do not make the assumption that parts must be watertight, we compute part occupancies using the same strategy as for our training losses in Section 3.2. For generation, we report Minimum Matching Distance (MMD) and Coverage Score (COV) (Achlioptas et al., 2018), using CD as the distance metric.
4.1 Shape and Part Reconstruction
3DS2VS Ours-1P DAE-Net BAE-Net PQ-Net Ours GT
@@ 4.1 Shape and Part Reconstruction
Figure 4: Shape reconstruction. Reconstruction of test shapes with all models. For part-based method, we color each part with a different color and translate the helix outside of the mixers for visualization. Best viewed digitally.
We evaluate the accuracy of shape and part reconstruction across all datasets and methods. To encode new test shapes at inference time, our model relies on auto-decoding: It remains frozen while reconstruction losses are minimized with respect to the latent vectors. We also report results for PartSDF with a single part, i.e., using our decoder in a part-agnostic manner, and refer to it as Ours-1P. We reconstruct meshes using Marching Cubes (Lorensen & Cline, 1987; Lewiner et al., 2003) at a resolution of $256$ for all methods. We report quantitative results in Table 1, with qualitative ones shown in Figure 4. Across all metrics, our method achieves the best results on cars and chairs, and equivalent results on mixers to Ours-1P, which does not reconstruct individual parts. We note that 3DShape2VecSet particularly struggles on this data, creating spurious floating surfaces and missing the central helices. Other part-based methods fall significantly behind on both surface and volume metrics.
@@
One thing that handicaps part-aware baselines is their reliance on voxelized data despite being INR-based: For speed and memory, the data is binarized and voxelized at $64^{3}$, which delivers efficient and fast models but greatly limits accuracy. This is particularly visible in the case of the thin helix of the mixers or the equally thin chair parts: They tend to either disappear or be inflated. Instead, PartSDF is trained with samples directly obtained from the original shape $\mathcal{S}$ and yields superior accuracy, capturing detailed part-specific geometry and structure. It is likely that approaches relying on more complex architectures (Deng et al., 2022; O. et al., 2022; Li et al., 2024) would be more competitive in reconstruction accuracy. However, as discussed in Section 2, they are not designed to provide composite shape priors usable for tasks such as part-based optimization.
@@
To illustrate that part encoder models can be used in conjunction with PartSDF, we also reconstruct the test set in a part agnostic manner: We encode each shape’s surface point cloud to obtain initial latents and poses, which are then refined with our auto-decoding strategy but without $\mathcal{L}_{\text{part}}$. As encoder, we use 3DShape2VecSet of reduced size, without the final cross-attention and, to produce an ordered set of latents and poses, learnable queries at the shape encoding level. As can be seen in the last row of Table 1, this allows us to accurately recover shapes and their parts, even though the thin structures of the mixers are more challenging.
@@ 4.2 Shape Generation
Table 2: Shape generation. We compute the Minimum Matching Distance MMD (×$\times 10^{4}$) and Coverage Score (COV) (%), using the Chamfer Distance (CD), between generated and test shapes.
@@
Ours 20.67 11.97 55.16 85.71 86.67 85.77
3DShape2VecSet
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
PQ-Net
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ours
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ours
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Shape generation. We generate shapes on all datasets and we also provide examples of pose conditioned generation (Ours) where part latents are generated based on the poses’ coarse description of the shape (left image for each pair). We translate the helix outside of the mixer for visualization.
We compare PartSDF’s shape generation abilities against those of 3DShape2VecSet and PQ-Net. The former relies on a diffusion model and the latter on a latentGAN (Achlioptas et al., 2018) to yield latent vectors from which shape or parts are decoded. For PartSDF, we leverage SALAD (Koo et al., 2023), a cascaded diffusion model that first generates part poses and then part latents conditioned on these poses. This also enables us to generate shapes fitting specified pose decompositions, something that PQ-Net cannot do with its single latent space. We report MMD and COV metrics in Table 2 and show generated examples in Figure 5. PartSDF achieves consistently better results than PQ-Net, with more detailed composite shapes. Our results are also better than, or on-par with, those 3DShape2VecSet, with the added bonus that our models are part-aware whereas those of 3DShape2VecSet are not, and therefore not usable for engineering design. Furthermore, it requires training different models for shape reconstruction and generation, while we use only one for both.

### v2

@@ 3.2 Cross-Part Auto-Decoder
Figure 4: Supervision for non-watertight parts. (a) Semantic parts of a chair, (b) a 2D slice of its signed distances (red/blue) with parts highlighted in color, and (c) the specific regions of space where each part is supervised. This enables training without requiring parts to be watertight.
3.3 Learning from Global Shape Supervision
Let us consider a dataset $\mathcal{D}$ where each shape $\mathcal{S}$ is decomposed into up to $P$ parts, given as the segmentation of its surface, as commonly found in online databases. To establish initial poses for each part, we fit simple primitives—cuboids or cylinders—to the segmented parts and use these primitives’ poses as the part poses. Therefore, each element of the dataset becomes a tuple $\mathcal{D}=\left\{\left(\mathcal{S},\left\{\mathcal{P}_{p}\right\},\left\{\mathbf{p}_{p}\right\}\right)_{i}\right\}$ of the shape, its parts, and their poses $\mathbf{p}_{p}=\left(\mathbf{q}_{p},\mathbf{t}_{p},\mathbf{s}_{p}\right)$.
@@ 3.3 Learning from Global Shape Supervision
For watertight shapes $\mathcal{S}$, PartSDF’s decoder is then trained in an auto-decoding fashion (Park et al., 2019), where model parameters $\theta$ and part latent vectors $\mathbf{z}_{p}$ are optimized jointly. Note that we can directly use the parts poses from $\mathcal{D}$ during training. When a part is missing from a specific shape, we assign it the average pose of that part over the dataset, allowing the model to naturally learn to output $\text{SDF}>0$ for nonexistent parts, without having to predict existence scores, as in Petrov et al. (2023).
Some parts $\mathcal{P}_{p}$, however, may not be watertight as they correspond to open surface fragments rather than closed volumes—a common characteristic of online part-based CAD models. This makes it difficult to compute reliable per-part signed distance fields, especially near part boundaries or in regions where parts overlap. While one option is to repair or approximate each part’s volume (Wu et al., 2020; Petrov et al., 2023), this often introduces artifacts and scaling issues. Instead, we take advantage of a key observation: when reconstructing the global shape as the minimum across all part SDFs, only the part whose surface is closest to a given point affects the output. We therefore supervise each part using the global shape’s SDF, but only in regions of space where that part is closest to the surface.
Formally, a part $\mathcal{P}_{p}$ is supervised only at points
$\mathbf{x}\in\left\{\mathbf{x}\in\mathbb{R}^{3}\,|\,p=\arg\min_{i}d_{i}(\mathbf{x})\right\},$ (5)
where $d_{i}(\mathbf{x})$ is the projection distance of $\mathbf{x}$ to part $i$, as depicted by Figure˜4. This region-based supervision ensures that each part learns only the portion of space it is responsible for and avoids penalizing it in regions where it is occluded or irrelevant. Crucially, it allows us to train directly from surface-based part annotations, without requiring watertight meshes or explicitly constructing per-part SDFs.
To further encourage modularity, we introduce a non-intersection loss that penalizes overlapping negative SDF predictions across parts. This promotes clean spatial separation, ensuring that each part occupies a distinct region of space—crucial for enabling localized editing, part replacement, and consistent composition under the minimum-SDF fusion.
@@
with an additional index $p_{i}$ of the part closest to $\mathbf{x}_{i}$, re-using $\mathcal{X}$ as a slight abuse of notation to explicitly state that we compute these losses on the same 3D samples. $\hat{s}_{p_{i}}$ is the predicted part SDF at those points.
Finally, we introduce a non-intersection loss that pushes the SDF of parts to be positive at each position $\mathbf{x}_{i}$ where at least two parts have SDF<0<0. Using $\hat{\mathcal{X}}$ as the subset of such 3D points, the loss is written
@@
where $\hat{\mathbf{s}}_{i}$ is the vector of predicted part SDF at point $\mathbf{x}_{i}$ and $\mathbf{w}_{i}=\text{softmax}\left(\tilde{\mathbf{s}}_{i}\right),$ defining $\tilde{\mathbf{s}}_{i}$ as $\hat{\mathbf{s}}_{i}$ with all positive values replaced by -$\infty$. This pushes towards >0>0 more strongly the SDFs that are closer to it.
@@ 3.4 Part Encoding and Generation
Once PartSDF’s decoder is trained, it is frozen to be used for downstream tasks, optionally in conjunction with separate specialized networks. While this requires training two networks independently, it makes the training of each one easier. Furthermore, the decoder need only be trained once and used again and again for the different tasks.
@@
To reconstruct a shape from a given modality, e.g., point clouds or images, encoders can be trained to directly predict part latents and poses, to then be decoded with PartSDF: With input data $\mathcal{I}$ corresponding to the shape of our training data $\mathcal{D}$, encoders are trained to map this new modality to the part latents and poses of our pre-trained decoder. As an example, we show how to reconstruct unseen shapes for which part decomposition is unknown in Section˜4.1. We do this by training a point cloud encoder to predict part latents and poses, which are then refined in an auto-decoding strategy and part-agnostic manner.
@@
For shape generation with a coherent part set, generative models can be trained to map random noise to the parts parametrization. Generative Adversarial Networks (GAN) (Goodfellow et al., 2014) have been used to generate shape latents, often dubbed latentGANs in related work (Achlioptas et al., 2018; Chen & Zhang, 2019) such as PQ-Net (Wu et al., 2020). Recently, diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) have gained a lot of traction for their generative performances, even in shape (Zhang et al., 2023) or part-based generation (Koo et al., 2023). Generating new shapes with PartSDF can be achieved by training such generative models on the part poses and latents of our pre-trained decoder and, even if not shown here, could be conditioned on images or texts (Zhang et al., 2023; Koo et al., 2023).
@@ 4 Experiments
We demonstrate our method on several tasks and compare it to multiple part-based baselines. Additional implementation and result details are in Appendices˜B and E.
Datasets
Public shape datasets like ShapeNet (Chang et al., 2015) and PartNet (Mo et al., 2019) often suffer from over-segmentation or inconsistent part definitions. To address this, we use three curated datasets with clean, consistent part decompositions: (1) Car, a hand-processed ShapeNet subset with separated wheels; (2) Mixer, liquid mixers with a helix, tube, and two attach points (Vasu et al., 2022); and (3) Chair, a cleaned PartNet-based set with individually segmented legs and arms. These datasets contain 1046, 1949, and 1332 shapes with 5, 4, and 8 parts, respectively. We use 80% for training and 20% for testing.
Baselines.
We compare against part-based methods across different supervision levels to evaluate their ability to represent composite shapes: DAE-Net (Chen et al., 2023) (unsupervised deformable parts), BAE-Net (Chen et al., 2019) (weakly supervised with 8 labeled shapes), PQ-Net (Wu et al., 2020) and PASTA (Li et al., 2024) (fully supervised, though PASTA does not output parts). We also evaluate 3DShape2VecSet (Zhang et al., 2023), a state-of-the-art, non-part-based method that, while unsuitable for composite shapes, provides an upper bound on achievable accuracy without enforcing part structure.
Metrics.
Reconstruction accuracy is assessed using three different metrics: Chamfer-Distance (CD) for surface accuracy, Intersection over Union (IoU) for volume, and Image Consistency (IC) (Guillard et al., 2022) for shape appearance and normals. Part reconstruction is evaluated by averaging the per-part IoU and as we do not make the assumption that parts must be watertight, we compute part occupancies using the same strategy as for our training losses in Section˜3.3. For generation, we report Minimum Matching Distance (MMD) and Coverage Score (COV) (Achlioptas et al., 2018), using CD as the distance metric.
4.1 Shape and Part Reconstruction
We evaluate the accuracy of shape and part reconstruction across all datasets and methods. At inference, our model uses auto-decoding: the decoder remains frozen while latent vectors are optimized to minimize reconstruction loss. We also report results for PartSDF using a single part and refer to it as Ours-1P. Meshes are reconstructed using Marching Cubes (Lewiner et al., 2003) at a resolution of $256$. We report quantitative results in Table˜1, with qualitative ones shown in Figure˜5. Across all metrics, our method achieves the best results on cars and chairs and equivalent results on mixers to Ours-1P, which does not reconstruct individual parts. Notably, 3DShape2VecSet struggles with the mixer’s helical structures, likely due to its positional encoding’s difficulty with periodic geometry. Other part-based methods perform significantly worse on both surface and volume metrics.
Table 1: Shape reconstruction. We compute the average Chamfer-Distance CD ($\times 10^{4}$), Intersection over Union IoU (%), and Image Consistency IC between reconstructions and corresponding test shapes. We also report average per-part Intersection over Union pIoU (%) for part-based methods. Ours-1P uses a single part to reconstruct the full shape and Ours-PC uses a point cloud encoder to get initial part latents and poses that are refined in a part-agnostic manner.
@@ 4.1 Shape and Part Reconstruction
PASTA 12.15 6.27 40.89 83.88 50.83 60.36 0.855 0.942 0.750 - - -
Ours 1.27 1.60 1.30 98.02 94.43 97.17 0.931 0.978 0.942 94.89 90.42 93.67
Ours-PC 1.28 1.73 1.57 97.95 83.38 97.09 0.930 0.966 0.942 94.25 78.61 88.13
3DS2VS Ours-1P DAE-Net BAE-Net PQ-Net PASTA Ours GT
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Figure 5: Shape reconstruction. Reconstruction of test shapes with all models. For part-based methods, we color each part with a different color and translate the helix outside of the mixers for visualization.
@@
One thing that handicaps DAE-Net, BAE-Net, and PQ-Net is their reliance on voxelized data despite being INR-based: For speed and memory, the data is binarized and voxelized at $64^{3}$, which delivers efficient and fast models but greatly limits accuracy. This is particularly visible in the case of the thin helix of the mixers or the equally thin chair parts: they tend to either disappear or be inflated. Instead, PartSDF is trained with samples directly obtained from the original shape $\mathcal{S}$ and yields superior accuracy, capturing detailed part-specific geometry and structure. While the more recent PASTA is also trained with these samples and improves on the other works, it still performs less well than our approach. Because it only uses the part bounding boxes as input, it may fail to capture the correct local details and, thus, fail to properly reconstruct some shapes, as in the chairs of Figure˜5. It is also unable to represent individual parts, despite being part-aware.
@@
Additionally, to illustrate that part encoder models can be used in conjunction with PartSDF, we also reconstruct test shapes using part-agnostic point clouds: We use a point cloud encoder based on a simplified 3DShape2VecSet without final cross-attention. It predicts initial part latents and poses, refined via auto-decoding without $\mathcal{L}_{\text{part}}$. As shown in the last row of Table˜1, this approach successfully recovers both global shape and parts, though very thin features remain more challenging.
@@ 4.2 Shape Generation
Table 2: Shape generation. We compute the Minimum Matching Distance MMD ($\times 10^{4}$) and Coverage Score (COV) (%), using the Chamfer Distance (CD), between generated and test shapes.
@@
PASTA 27.72 16.31 118.35 40.95 51.03 23.97
Ours 20.67 11.97 55.46 85.71 86.67 83.90
3DShape2VecSet PQ-Net PASTA Ours Ours
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Shape generation. We generate shapes on all datasets, and we also provide examples of pose-conditioned generation (Ours) where part latents are generated based on the poses’ coarse description of the shape (left image for each pair). When possible, we translate the helix outside of the mixer for visualization.
We compare PartSDF’s shape generation abilities against those of 3DShape2VecSet, which relies on a diffusion model to generate its set of latent vectors, PQ-Net, which trains a latentGAN (Achlioptas et al., 2018) to yield a global latent vector, and PASTA, which uses an autoregressive transformer to generate part bounding boxes. For PartSDF, we leverage SALAD (Koo et al., 2023), a cascaded diffusion model that first generates part poses and then part latents conditioned on these poses. This also enables us to generate shapes fitting specified pose decompositions, something that PQ-Net cannot do with its single latent space, and to generate various geometries for a given pose decomposition, which PASTA cannot do as it only relies on part poses for reconstruction. We report MMD and COV metrics in Table˜2 and show generated examples in Figure˜6. PartSDF achieves consistently better results than the part-aware baselines, with more detailed composite shapes. We note that despite significant integration effort on our side, the PASTA-generated poses exhibit low diversity and occasional pose errors. Our results are also better than, or on par with, those of 3DShape2VecSet, in addition that our models are part-aware, whereas those of 3DShape2VecSet are not, and therefore not usable for engineering design. Furthermore, it requires training different auto-encoder models between shape reconstruction and generation, while PartSDF uses the same decoder for both.
4.3 Part Manipulation